# DensiCrafter: Physically-Constrained Generation and Fabrication of Self-Supporting Hollow Structures

### v1

@@ Abstract
Codehttps://github.com/dangsq/DensiCrafter
@@ Introduction
$\mathbf{c}=\frac{\displaystyle\int_{\Omega}\mathbf{x}\,\tilde{\rho}(\mathbf{x})\,\mathrm{d}\mathbf{x}}{\displaystyle\int_{\Omega}\tilde{\rho}(\mathbf{x})\,\mathrm{d}\mathbf{x}}$ (7)
where $z_{\min}$ means the ground‐level of $O$.
@@
$\mathcal{L}_{\mathrm{center}}=\frac{1}{|S|}\int_{\mathbf{s}\in S}\bigl\lVert(s_{x},s_{y})-(c_{x},c_{y})\bigr\rVert_{2}\,\mathrm{d}\mathbf{s}$ (8)
@@
$\mathcal{L}_{\mathrm{region}}=-\,|S|=-\int_{\mathbf{s}\in S}\mathrm{d}\mathbf{s}$ (9)
Height Loss. Lowering the center of mass increases the critical overturning angle, thereby enhancing stability against rotational perturbations. To encourage this, we penalize $c_{z}$:
@@
To minimize the object’s total mass while preserving its overall shape, we introduce two regularizations: the mass penalty and the strategically restricted optimization domain.
@@
$\mathcal{L}_{\mathrm{mass}}=\int_{\Omega}\tilde{\rho}(\mathbf{x})\,\mathrm{d}\mathbf{x}$ (11)

### v2

@@ Abstract
Codehttps://github.com/idvxlab/DensiCrafter
@@ Introduction
$\mathbf{c}=\frac{\displaystyle\int_{\Omega}\mathbf{x}\,\tilde{\rho}(\mathbf{x})\,\mathrm{d}V}{\displaystyle\int_{\Omega}\tilde{\rho}(\mathbf{x})\,\mathrm{d}V}$ (7)
where $z_{\min}$ means the ground‐level of $O$ and $\mathrm{d}V$ denotes the differential volume element.
@@
$\mathcal{L}_{\mathrm{center}}=\frac{1}{|S|}\int_{\mathbf{s}\in S}\bigl\lVert(s_{x},s_{y})-(c_{x},c_{y})\bigr\rVert_{2}\,\mathrm{d}A$ (8)
where $\mathrm{d}A$ denotes the differential area element.
@@
$\mathcal{L}_{\mathrm{region}}=-\,|S|=-\int_{\mathbf{s}\in S}\mathrm{d}A$ (9)
Height Loss. Lowering the center of mass increases the critical overturning angle, thereby enhancing stability against rotational perturbations. Therefore, we penalize $c_{z}$:
@@
To minimize the total mass while preserving the overall shape, we introduce two regularizations: the mass penalty and the restricted optimization domain.
@@
$\mathcal{L}_{\mathrm{mass}}=\int_{\Omega}\tilde{\rho}(\mathbf{x})\,\mathrm{d}V$ (11)